With the proper figures, describe the following terms in the context of Series R-L-C circuit

(i) Resonance Frequency (f0)

(ii) Power factor for f = f0,

(iii) Half-Power Bandwidth

(iv) Selectivity (Quality Factor) 


1
Expert's answer
2022-03-13T18:51:43-0400

RLC circuit

(1) resonance frequency

"X_L=X_C"

"wL=\frac{1}{wc}"


"w=\frac{1}{\sqrt{LC}}"

"f=\frac{1}{2\pi\sqrt{LC}}"

(B)

Power factor

"Cos\theta=\frac{R}{Z}"

"Z=\sqrt{R^2+(X_L-X_C)^2}"

"(c)"

Band width

"BW=f_2-f_1=\frac{R}{2\pi L}"

(D)

Quality factor

"Q=\frac{w_rL}{R}=\frac{1}{w_rCR}"